A 1.00kg ball travelling towards a small soccer player at a velocity of 6.00m/s rebounds off the soccer players foot at a velocity of 7.00m/s. If the time of contact between the ball and the players foot was 3.00 × 10² seconds, what was the force that the foot applied on the ball?


1
Expert's answer
2021-04-26T16:59:17-0400
"F=m\frac{v-u}{t}\\F=(1)\frac{7-(-6)}{0.03}=433\ N"